# Weakly Supervised Gaussian Contrastive Grounding with Large Multimodal Models for Video Question Answering

### v1

@@ 5.3 Results and Analysis
The experimental results in Table 1 and 2 show the superiority of our weakly-supervised GCG compared to existing approaches. For the multi-choice setting, we achieve an accuracy of 74.6%, 72.1%, and 73.1% in NExT-QA, Causal-VidQA, and Intent-QA respectively. For the open-ended setting, we achieve an accuracy of 61.7%, 49.5%, and 49.9% in MSVD-QA, MSRVTT-QA, and ActivityNet-QA respectively. To ensure a fair comparison, we also apply the same settings to get the results for the vanilla InstructBLIP on these datasets as baselines, with $K=4$ frames uniformly sampled from the $T=32$ frames as visual inputs. Although the baseline InstructBLIP performed fair on these datasets, our proposed GCG showed a significant improvement, particularly in questions that require complex causal-temporal reasoning (+2.1% and +2.7% for Tem and Cau in NExT-QA, +2.3% for Cou in Causal-VidQA). Moreover, despite not having undergone video-text pretraining, our method still surpasses those pre-trained VideoQA models (e.g., HiTea, VALOR) on MSVD-QA, MSRVTT-QA, and ActivityNet-QA. We also observe that the improvements on ActivityNet-QA (+3.6%) are larger than the MSVD-QA (+2%) and MSRVTT-QA (+2.5%). This can be attributed to the average video length of ActivityNet-QA being 180 seconds, which is much longer than MSVD-QA (10s) and MSRVTT-QA (15s), demonstrating the necessity of discovering question-critical moments.
@@
Table 4: Ablation studies on loss components of GCG.
Figure 4: Qualitative results on NExT-QA test set. The frames selected by our method are highlighted in blue dashed lines. The ground truth answers are in green. We also display the weakly grounded similarity scores of each frame.
Figure 5: Ablation studies on hyperparameters $T$ and $\sigma$.
5.4 Pseudo-labels Analysis.
We further explore the performance with different frames as visual inputs to verify the pseudo-labeled $w$: Pseudo-labeled (E) means we choose the $K$ frames whose indexes correspond to $w\in\mathbb{R}^{K}$ as visual inputs (detailed in Section 4.2). Pseudo-labeled (Q) is obtained the same way as Pseudo-labeled (E) but with the pure question for similarity computation. Baseline means we uniformly sample $K$ frames as visual inputs.
Table 3 shows that Pseudo-labeled (E) exhibits significantly improved performance, particularly in benchmarks featuring longer videos and more complicated questions like NExT-QA and Intent-QA. This verifies the effectiveness of using the event description to provide pseudo-labels as weak supervision. Moreover, Pseudo-labeled (E) performs much better than Pseudo-labeled (Q). This can be explained from two perspectives: (1) The event descriptions include the contents of the answers needed to be grounded, filling the semantic gap between the pure question and the answer. (2) The CLIP models are mostly pre-trained on images and declarative texts, therefore the declarative event descriptions are more suitable for similarity computation to decide keyframes.
5.5 Ablation Studies
We investigate the role of our weakly-supervised framework with different variants and hyperparameters of GCG.
@@ 5.5 Ablation Studies
  • $\mathcal{L}_{vqa}+\mathcal{L}_{reg}$ and $\mathcal{L}_{vqa}+\mathcal{L}_{con}$ match equally in accuracy that consistently surpasses baseline and $\mathcal{L}_{vqa}$. $\mathcal{L}_{reg}$ is responsible for regularizing the indexes of peaks in $p$ to approximate the timestamps of weakly grounded $w$, and $\mathcal{L}_{con}$ ensures the maximization between the selected moments and event descriptions.
@@
5.6 Qualitative results.
@@
We also present qualitative results in Figure 4 to gain a clearer insight into our method, along with the frames identified by GCG (bounded with blue dashed lines) and the weakly grounded similarities computed by the event description. Both cases show semantic correspondence between the question and the selected moments. For instance, in the upper case, the frames identified by GCG precisely correspond to the event [The baby rolled the roller on the toy when she reached it] with relatively higher similarity scores, demonstrating the strong ability of our GCG to localize video moments crucial for answering the question. By leveraging the information from this localized segment, the LMM can successfully arrive at the correct answer [ roll the roller on the toy]. In contrast, without our GCG, the presence of massive redundancy in the video overwhelms the reasoning process and leads to a false prediction of [laugh loudly].
6 Conclusion
@@
In this paper, we have studied the problem of discovering question-critical moments in videos when adapting LMMs for VideoQA tasks. To address the shortcomings of the uniform sampling strategy and the absence of human annotations for question-critical timestamps in VideoQA datasets, we introduce a weakly-supervised framework to force the LMMs to reason out the answers by grounding question-critical moments. We propose the Gaussian-based Contrastive Grounding, a flexible and lightweight method to dynamically select keyframes with end-to-end training. Through a series of experiments and analyses, we have demonstrated the effectiveness of our approach in various challenging VideoQA tasks, particularly excelling in causal-temporal reasoning. For future work, we intend to focus on mitigating the biases and superficial correlations observed in previous studies, and further enhance the reasoning abilities of current models.

### v3

@@ 5.4. Main Results
Table 1 and 2 show the superiority of our method. For the multi-choice setting, we achieve an accuracy of 74.6%, 72.1%, and 73.1% in NExT-QA, Causal-VidQA, and Intent-QA respectively. For the open-ended setting, we achieve an accuracy of 61.7%, 49.5%, and 49.9% in MSVD-QA, MSRVTT-QA, and ActivityNet-QA respectively. To ensure a fair comparison, we also apply the same settings to get the results for the vanilla InstructBLIP on these datasets as baselines, with $K=4$ frames uniformly sampled from the $T=32$ frames as visual inputs.
@@
Table 3. Ablation studies on loss components of GCG.
Figure 5. Ablation studies on hyperparameters $T$ and $\sigma$.
Although the baseline InstructBLIP performed fair on these datasets, our proposed GCG showed a significant improvement, particularly in questions that require complex causal-temporal reasoning (+2.1% and +2.7% for Tem and Cau in NExT-QA, +2.3% for Cou in Causal-VidQA). When both using BLIP-2 as the baseline model, GCG still performs better than SeViLA (Yu et al., 2023) (73.6% vs 71.5% in NExT-QA) without extra pre-training or a multi-stage training scheme. Moreover, despite not having undergone video-text pretraining, our method still surpasses those large-scale pre-trained models (e.g., HiTea, COSA, VALOR) on MSVD-QA, MSRVTT-QA, which primarily feature straightforward questions and short videos (10-15s). We also observe that the improvements on ActivityNet-QA (+3.6%) are much larger than the MSVD-QA (+2%) and MSRVTT-QA (+2.5%). This can be attributed to the average video length of ActivityNet-QA being 180 seconds, which is much longer than MSVD-QA (10s) and MSRVTT-QA (15s), emphasizing the necessity of discovering question-critical moments more with our method.
5.5. Ablation Studies
We investigate the role of our framework with different variants and hyperparameters of GCG, by using InstructBLIP as the baseline model, and NExT-QA and MSVD-QA as the default benchmarks.
@@ 5.5. Ablation Studies
  • $\mathcal{L}_{vqa}+\mathcal{L}_{reg}$ and $\mathcal{L}_{vqa}+\mathcal{L}_{con}$ match equally in accuracy that consistently surpasses baseline and $\mathcal{L}_{vqa}$. $\mathcal{L}_{reg}$ is responsible for regularizing the indexes of peaks in $p$ to approximate the timestamps of weakly grounded $w$,
    Figure 6. Ablation studies on the number of negative samples.
    Layer Num. $N$ Hidden Size $D_{G}$ Params. NExT-QA (%)
    2 1024 27.6M 74.8
    2 768 15.8M 74.1
    2 512 7.3M 74.3
    2 256 2.0M 74.6
    1 256 1.2M 73.9
    2 256 2.0M 74.6
    3 256 2.8M 74.4
    4 256 3.6M 74.4
    Table 4. Ablation studies on the Gaussian generator.
    Datasets Baselines +GCG +GCG && LoRA
    NExT-QA 72.5 74.6 75.5
    Causal-VidQA 70.6 72.1 73.4
    Intent-QA 71.5 73.1 74.5
    MSVD-QA 59.7 61.7 62.2
    MSRVTT-QA 47.0 49.5 49.8
    ActivityNet-QA 46.3 49.9 50.1
    Table 5. InstructBLIP with both GCG and LoRA.
    and $\mathcal{L}_{con}$ ensures the maximization between the selected moments and event descriptions.
@@
Figure 7. Qualitative results on NExT-QA test set. The frames selected by our method are highlighted in blue dashed lines. The ground truth answers are in green. We also display the weakly grounded similarity scores of each frame.
Ablations on negative sample numbers $N_{intra}$ and $N_{inter}$. We also investigate how the number of intra-negative samples $N_{intra}$ and inter-negative samples $N_{inter}$ influence the performance. As verified in Figure 6 (a), sampling enough intra-negative moments is beneficial to mining the positive moments for reasoning. Figure 6 (b) shows that the performance gains also increase as the number of inter-video negative moments increases. Moreover, the impact of intra-negative moments within the same video is larger than inter-negative moments, because the scenes in the same video are more confusing to distinguish the true question-critical moments. We also observe that the performance decreases after the $N_{intra}$ and $N_{inter}$ reach a specific value of 16 and 32 respectively. We argue that selecting excessive negative moments tends to distract positive moments and therefore degrades the performance. We finally set the $N_{intra}=16$ and $N_{inter}=32$ which balance well the performance and computational cost.
Ablations on hidden size $D_{R}$ and layer number $N$. To determine the optimal configuration for our Gaussian generator, we set different layer number $N$ and hidden size $D_{R}$ for the transformer encoder in the Gaussian generator. The results of these variants on NExT-QA are shown in Table 4. We can see that when the layer number of the encoder is fixed to $N=2$, the model achieves the best performance of $74.8\%$ with $D_{G}=1024$. However, such improvement in performance comes at the cost of a significant increase in the number of parameters (27.6M). To ensure the flexibility and adaptability of our method, we choose $D_{G}=256$ as our default setting, which achieves a balance between fair performance (74.6%) and much fewer parameter quantities (2.0M). We can also observe that when the hidden size is fixed to $D_{G}=256$ and the layer number $N\geq 2$, the change in $N$ has a relatively minor impact on performance. For better computation efficiency, we adopt $N=2$ for our Gaussian generator.
Further LoRA tuning to push better results. We note that LMMs with GCG can be considered a strong model for VideoQA. To achieve better performance, we further add per-task LoRA tuning (Hu et al., 2021) for the frozen language model in LMMs (using a rank of 16), yielding new SOTA results on these VideoQA benchmarks. These new SOTA results in Table 5 indicate the extensibility and flexibility of the design of our GCG, which can be easily combined with other components for LMMs.
5.6. Qualitative results.
@@
We also present qualitative results in Figure 7, along with the frames identified by GCG (bounded with blue dashed lines) and the weakly grounded similarities scores. Both cases show semantic correspondence between the question and the selected moments, enhancing the interpretability of LMMs for VideoQA by revealing which visual scenes result in the answers. In the upper case, the frames identified by GCG precisely correspond to the event [The baby rolled the roller on the toy when she reached it] with relatively higher similarity scores, demonstrating the ability of GCG to localize video moments crucial for answering the question. By leveraging the information from this localized segment, the LMM can successfully arrive at the correct answer [roll the roller on the toy]. In contrast, without our GCG, the presence of massive redundancy in the video overwhelms the reasoning process and leads to a false prediction of [laugh loudly].
6. Conclusion
@@
In this paper, we have studied the problem of discovering question-critical moments in videos when adapting LMMs for VideoQA tasks. To address the shortcomings of the uniform sampling strategy and the absence of human annotations for question-critical timestamps in VideoQA datasets, we introduce a weakly-supervised framework to force the LMMs to reason out the answers by grounding question-critical moments as visual inputs. To achieve this, we utilize the CLIP models to automatically provide the pseudo-labeled timestamps of keyframes. With these keyframes as additional weak supervision, we propose the Gaussian-based Contrastive Grounding, a flexible and lightweight method to dynamically select question-critical moments with end-to-end training. Through a series of experiments and analyses, we have demonstrated the effectiveness of our approach in various challenging VideoQA tasks, particularly excelling in causal-temporal reasoning.
Limitations. Despite the significantly improved performance on several VideoQA datasets with our method for LMMs, it’s essential to acknowledge the potential presence of language bias in the frozen language models. In our future work, we plan to mitigate these biases and enhance the reasoning ability of current LMMs further.